Let A and B be two incompatible events such that P (A) = 0, 3 and P (B) = 0, 5. What are
the probabilities
(a) A or B is realized;
b) that A, but not B, is realized;
c) that neither A nor B are realized?
Answer #1

a) A and B are incompatible events i.e., they are mutually exclusive

Therefore, the probability that either A or B is realized = p(A)+p(B) - p(AB)

Since they are mutually exclusive, p(AB) = 0

Therefore, p(A)+p(B) = 0.8

b) A but not B is realized = P(A)+p(B|) = 0.3+(1-0.5) = 0.8

c) Neither A nor B = 1-p(either A or B) = 1-0.8 = 0.2